# TextFusion: Unveiling the Power of Textual Semantics for Controllable Image Fusion

### 21 Dec 2023 v1

@@ 4.5 Ablation study relating to the affine fusion unit
As shown in Table. 3, the image fusion assessments are also consistent with our observation of the qualitative results. In particular, using the proposed setting enables our results to obtain the best performance regarding the performance of preserving the texture details (Qabf) and maintaining the information fidelity (VIF). Besides, in terms of controllability and the imaging quality, our fusion strategy also has better performance than the tested alternatives.
4.6 Model size and performance comparison
5 Conclusion
6 Acknowledgement
This work was supported by the National Natural Science Foundation of China (62020106012, U1836218, 62106089, 62202205), the 111 Project of Ministry of Education of China (B12018), the Engineering and Physical Sciences Research Council (EPSRC) (EP/R018456/1, EP/V002856/1), and the Postgraduate Research & Practice Innovation Program of Jiangsu Province (KYCX23_2525).
Supplementary Material
7 More details about the IVT dataset
Figure 12: An illustration of the training process of the proposed model. In our implementation, we assign different levels of salience for the interested and irrelevant regions, based on activity level maps (pixel intensity) and information measurement (edge gradient), to obtain the weights for the input images. Accordingly, a fusion network is optimized by using a loss function reflecting the salience differences, with the input text information used to guide the fusion process.
8 More details about the loss function
As we mentioned in Section 3, our loss function is formulated as
$Loss_{\mathrm{total}}=Loss_{\mathrm{int}}+Loss_{\mathrm{irr}},$ (9)
where $Loss_{\mathrm{int}}$ and $Loss_{\mathrm{irr}}$ stand for the region of interest and the irrelevant region loss items, respectively.
In Fig. 12, we provide an illustration of the loss function. Firstly, in the region of interest, we use the spatial attention strategy adopted in [43] to preserve the salient information from the infrared and visible images in terms of pixel intensity. However, different from these methods, we expect to realize the fusion process in an end-to-end manner. To achieve this objective, we adopt the pre-trained autoencoder network [16] to extract the deep image features and use $\phi_{i}(I)$ to denote the $i$-th channel of the feature maps. Subsequently, we conduct the $l_{1}$ normalisation on these features to calculate the activity level map
$A(I)(x,y)=\|\phi_{1:C}(I)(x,y)\|_{1},$ (10)
where A denotes the activity level map and $C$ indicates the number of channels. The activity level map can quantify the salient content of input images [44]. Since information with a higher response value is favoured by this strategy, the region of interest can be well highlighted accordingly. Following this, we conduct a softmax operation on $A(I_{\mathrm{ir}})$ and $A(I_{\mathrm{vis}})$ to obtain the weight maps
$[w_{\mathrm{ir}},w_{\mathrm{vis}}]=\mathrm{softmax}(A(I_{\mathrm{ir}}),A(I_{\mathrm{vis}})),$ (11)
where $w_{\mathrm{{ir}}}$, $w_{\mathrm{{vis}}}$ are the pixel-level weight maps of source images.
Next, we apply these two weights in the loss function to dynamically control the proportion of pixel information transferred from the input images into the fusion result. Thus, we obtain the formulation of this loss function
where $\|\cdot\|_{F}$ denotes the Frobenius norm.
On the other hand, for the irrelevant regions, the information measurement approaches based on the gradient information [2, 39] have been shown to handle the visual texture details well. We tend to retain this information in the irrelevant regions as supplementary for the salient interest part. Since we maintain significant contrast for the fused images, such differences in design can also benefit the downstream vision task to better index the position of the foreground objects.
Specifically, we use a similar gradient-based strategy to control the mixing of the source images in the non-salient areas. Firstly, we estimate the information content of input images using a pre-trained VGG-19 network [31]. This measurement is defined as
where $\theta_{i}^{j}(I)$ denotes the $j$-th channel of the feature maps obtained before the $i$-th max-pooling layer of the VGG model, $n_{i}$ is the number of the feature maps in the corresponding convolutional layer, $\nabla$ is the gradient operator and $\overline{X}$ is defined as the reverse of $X$, $i.e.$, $\overline{X}(x,y)=1-X(x,y)$. $DS(X,s)$ denotes the down-sampling operation with the scale factor of $s$ on matrix $X$.
We apply a softmax operation to this information measurement to obtain the mixing weights $p_{{ir}}$ and $p_{{vis}}$:
$[p_{\mathrm{ir}},p_{\mathrm{vis}}]=\mathrm{softmax}(IM(I_{\mathrm{ir}}),IM(I_{\mathrm{vis}})).$ (16)
Finally, the irrelevant region loss is formulated as follows:
8.1 More ablation experiments
In this section, we present more ablation experiments and visualisation results to highlight the importance of various components in our TextFusion approach.
@@ 8.1 More ablation experiments
8.1.1 Visualisation of the proposed affine fusion unit
@@
Table 4: Quantitative results of the ablation study about the network architecture. (D denotes the number of the transformer blocks)
@@
8.1.2 The impact of network depth $D$ in the vision encoder
In our TextFusion network, we use several Swin Transformer Blocks to build the vision encoders. In this section, we conduct ablation experiments to evaluate and select the optimal network architecture. Specifically, our TextFusion network is trained based on different settings of $D$ from the vision encoder part. As shown in Table. 4, with the increase of the network depth (volume), the fusion results tend to have better performance on the metrics of SSIM. However, although the similarity between the input and fused images is improved, the no-reference image metrics (SF and SD) reveal that the quality and clarity of the fusion results are not promising. In contrast, the fusion results obtained by solely utilizing one transformer block have better performance on these perspectives. This phenomenon can be attributed to the fact that the combination of higher-level semantics conveyed through the textual input is adequate for the low-level appearance image features to produce robust fused images.
8.1.3 The impact of the threshold $\alpha$ of the overlapping ratio
@@
8.1.4 Effectiveness of the loss function / textual guidance

### 08 Feb 2024 v2

@@ V-E1 Ablation study relating to the affine fusion unit
As shown in Table. III, the image fusion assessments are also consistent with our observation of the qualitative results. In particular, using the proposed setting enables our results to obtain the best performance regarding the ability of preserving the texture details (Qabf) and maintaining the information fidelity (VIF). Besides, in terms of controllability and the imaging quality, our fusion strategy also has better performance than the tested alternatives. It is worth noting that, when the visible modality is used to generate the weight term. With moderate controllability, the average performance of setting (c) on this dataset cannot catch that of setting (b). This may due to the reason that such control manner is not consistent with our loss function design. Consequently, it dose not well preserve salient thermal information as our TextFusion. Furthermore, using the average operation, the original metric item (average style) in our textual assessment also contributes to the higher performance of setting (b).
V-E2 The impact of the threshold $\alpha$ of the overlapping ratio
@@
V-E3 Visualisation of the proposed affine fusion unit
@@
TABLE IV: Quantitative results of the ablation study about the network architecture. (D denotes the number of the transformer blocks)
@@
V-E4 The impact of network depth $D$ in the vision encoder
In our TextFusion network, we use several Swin Transformer Blocks to build the vision encoders. In this section, we conduct ablation experiments to evaluate and select the optimal network architecture. Specifically, our TextFusion network is trained based on different settings of $D$ from the vision encoder part. As shown in Table. IV, with the increase of the network depth (volume), the fusion results tend to have better performance on the metrics of SSIM. However, although the similarity between the input and fused images is improved, the no-reference image metrics (SF and SD) reveal that the quality and clarity of the fusion results are not promising. In contrast, the fusion results obtained by solely utilizing one transformer block have better performance on these perspectives. This phenomenon can be attributed to the fact that the combination of higher-level semantics conveyed through the textual input is adequate for the low-level appearance image features to produce robust fused images.
@@ V-E4 The impact of network depth $D$ in the vision encoder
V-E5 Effectiveness of the loss function / textual guidance
@@
V-F Model size and performance comparison
VI Conclusion
Although our TextFusion has obtained promising results, this work still has some limitations. First of all, one of our contributions is the proposal of a new dataset, which consists of infrared and visible image pairs with textual descriptions. However, there is still a gap between the size of our dataset and existing benchmarks for other vision tasks. We will continue enlarging this dataset in the future. Secondly, in our implementation, we use the association maps to link each word and corresponding image patches. This straightforward mechanism cannot always locate the region of interest, which may impede the fusion performance. Lastly, our revised image fusion indexes demonstrate a greater consistency with the visualisation of different fusion results. But the rationale behind incorporating the fusion guidance into the evaluation metrics has not been fully explored. Research and discussion on quality metrics for image fusion is still an active and ongoing area of study.
VII Acknowledgement
This work was supported by the National Natural Science Foundation of China (62020106012, U1836218, 62106089, 62202205), the 111 Project of Ministry of Education of China (B12018), the Engineering and Physical Sciences Research Council (EPSRC) (EP/R018456/1, EP/V002856/1), China Scholarship Council (CSC), and the Postgraduate Research & Practice Innovation Program of Jiangsu Province (KYCX23_2525).